OpenStudy (anonymous):

P(t) = 1000+360sin((π/6)t)+180sin((π/3)t) Find the derivative and then solve for when P(t)=0

OpenStudy (amistre64):

its alot of sums; so deriving is easy

OpenStudy (amistre64):

0 +360pi/6 cos(t.pi/6) +180pi/3cos(t.pi/3)

OpenStudy (anonymous):

ok i got that part right. now the solving for p(t)=0 part is where im stuck

OpenStudy (amistre64):

looks tricky ;)

OpenStudy (amistre64):

360/6 = 180/3 so that all the same

OpenStudy (amistre64):

cos(t.pi/6) + cos(t.pi/3) = 0 when.....

OpenStudy (amistre64):

you agree with this so far?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

we gotta get them at 180 from each other than right?

OpenStudy (amistre64):

t.60 = t.30 + 180; or, t.30 = t.60 + 180

OpenStudy (amistre64):

t(60-30) = 180 t(30) = 180 t = 60 maybe?

OpenStudy (amistre64):

err...6 lol

OpenStudy (amistre64):

6.60 = 360 6.30 = 180; thats a good one right?

OpenStudy (amistre64):

a(1) + a(-1) = 0 right

OpenStudy (amistre64):

so every multiple of 6 should be good for that one...

OpenStudy (amistre64):

t.30 = t.60 + 180 t(30-60) = 180 t = 180/-30 t = -6; same thing

OpenStudy (amistre64):

so whenever t = a multiple of 6, we should be good

OpenStudy (anonymous):

i kinda lost you at t.60 = t.30 + 180; or, t.30 = t.60 + 180. from cos(t.pi/6) + cos(t.pi/3) = 0 in fact, for that^ i simplified and got 60cos((t x pi/6)+ (t x 2pi/6) = 0

OpenStudy (amistre64):

given that t not equal zero perhaps; or some such restriction

OpenStudy (amistre64):

there are bound to be multples of 6 that make t.60 = t.30; in that case its of no use

OpenStudy (amistre64):

ok; we need angles that are 180 from each other in order for there cosines to be opposites of each other right?

OpenStudy (anonymous):

uh huh

OpenStudy (amistre64):

so we start at 60 and 30 and need to at least find the angles that are 180 from each other.... t.60 = t.30 + 180 means that for every multiple of t, we move by 60 and 30 till we find angles that are 180 from each other right?

OpenStudy (amistre64):

t.60 - t.30 = 180 t(60 - 30) = 180 t = 180/30 t = 6

OpenStudy (anonymous):

ooh ok

OpenStudy (amistre64):

now when the angels are 360 from eah other they are the same...so we exclude doing the same thing but with +360 instead of +180 right?

OpenStudy (amistre64):

360/30 = 12; so every multiple of 12 wont work.... right?

OpenStudy (anonymous):

uh huh

OpenStudy (amistre64):

so odd multiples of 6 work; but even multiples fail; so we restricct it to when t is an odd multiple of 6

OpenStudy (amistre64):

i can also think of one more condition, but havent found a way to describe it yet......... it when they angle to the right matches the angle on the left; like a 60 and a 120 would have opposite cosines

OpenStudy (anonymous):

lol i think im good. thank you.

OpenStudy (amistre64):

:) ok

OpenStudy (anonymous):

another question if you dont mind? lol

OpenStudy (amistre64):

i dunno, im pretty much good for one question and then go stupid lol

OpenStudy (amistre64):

whats the q

OpenStudy (anonymous):

LOL, it's a position, velocity, acceleration question.

OpenStudy (anonymous):

Find the position function X(t), of a particle moving along a straight line, with constant acceleration equal to 2, knowing that the position and velocity vanish at t = 2.

OpenStudy (amistre64):

is it posted already or you wanna hit me with it

OpenStudy (anonymous):

Find the initial position and the time t > 0 at which the particle passes again through the initial position.

OpenStudy (anonymous):

i know i should do anti derivative for the first part in order to find the velocity equation.

OpenStudy (amistre64):

acceleration is the derivative of velocity is the derivative of position

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

with a constant acceleration of 2; that means we got linear velocity of 2x+c right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

would i plug in velocity = 0 since it vanishes or no?

OpenStudy (amistre64):

and a position of x^2 +bx +cthen

OpenStudy (amistre64):

the only way for a linear function to vanish at some point it to make a hole iin the graph right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

which means making the velocity function an extention of a rational function

OpenStudy (amistre64):

2x +c ------ at best d-2

OpenStudy (amistre64):

but with a d-2 up top

OpenStudy (anonymous):

what's d-2?

OpenStudy (amistre64):

something to make a whole at 2 to make the graph vanish

OpenStudy (anonymous):

lol i think you might be over thinking the problem. this was suppose to be an easy math problem my prof gave on the test. haha

OpenStudy (amistre64):

i could be over thinking it tho :)

OpenStudy (anonymous):

lmao would it be wrong to set the velocity - 0?

OpenStudy (anonymous):

velocity = 0*

OpenStudy (amistre64):

cant have a constant acceleration whout velcoty can we?

OpenStudy (amistre64):

my keyboard hates me lol

OpenStudy (anonymous):

haha its okay

OpenStudy (anonymous):

there was a similar question to this, and my sis helped me with that.

OpenStudy (anonymous):

and in that other problem, my sis taught me to set velocity = 0

OpenStudy (anonymous):

then solve for the constant

OpenStudy (amistre64):

cant say that i ever came across the problem before, so it might be plausible :)

OpenStudy (anonymous):

and then from that, do another anti derivative to find the position fuction

OpenStudy (anonymous):

lol alrighty, thanks anyways! enjoy the rest of your day :]

OpenStudy (amistre64):

:) yw

OpenStudy (toxicsugar22):

amristre can u help me